# Beyond Syntax: Action Semantics Learning for App Agents

### v1

@@ 3.3 Semantic Estimator
where $o_{t}$ is the textual description of the UI state $s_{t}$, $a_{t}$ is the textual description of an action, and $\Delta\hat{o}_{t}$ is the description of the impact of the action $a_{t}$ on the UI state $s_{t}$. In practice, we implement this world model based on the closed LLM API by leveraging the LLM’s strong capabilities in language understanding and world knowledge to simulate how the environment would evolve in response to the action. The prompt of our world model can be found in the Appendix.
Semantic calculator. With the state transitions generated by the world model, SEE employs a semantic calculator to generate the semantic reward for our ASL framework-generated transitions in two steps. Firstly, the semantic calculator uses BERT devlin2019bert to project the generated textual description for state transitions into a feature space. This process can be formulated as:
@@
This similarity score is used in our ASL framework as the semantic reward. A higher similarity score indicates a higher probability that a semantic equivalence exists between the action generated by the App agent and the ground truth action, providing a reward signal that guides the training of the agent towards semantically meaningful actions, even when the syntax may differ.
3.4 Theoretical Analysis
In this section, we demonstrate that enabling the App agent to learn action semantics by minimising the semantic loss function defined in Eq. (3) leads to an agent that is more robust to the OOD problems compared to an agent that purely optimises the learning objective formulated in Eq. (1). Mathematically, we present Theorem 3.2 with the proof in the Appendix.
Theorem 3.2.
4 Experiments
In this section, we first introduce our experimental setup. We then demonstrate the practical value of the proposed ASL framework on two online smartphone operation Apps benchmarks. To evaluate the robustness of the agent fine-tuned by our ASL framework, we use an offline smartphone operation benchmark that includes a clear separation between in-distribution and OOD data. Finally, we conduct an ablation study to study the contribution of components within the ASL framework.
4.1 Experimental Setup
Datasets. We conducted our experiments on three benchmark datasets: AndroidLab xu2024androidlab , AndroidWorld rawles2024androidworld , and AndroidControl android_control . AndroidWorld includes a high-fidelity simulator that emulates realistic Android device behaviour, enabling interactions with 20 real-world applications across 116 user-defined tasks. To determine whether a task is successfully completed, each task is paired with manually authored rules that provide precise success criteria. AndroidLab includes 138 predefined tasks across nine Android applications built on virtual devices. Each task is divided into multiple required page states as sub-goals, with UI tree structure matching verifying correct traversal. The AndroidLab environment ensures reproducibility and eliminates external network or time dependencies. AndroidControl comprises 15,283 human demonstrations spanning 14,548 unique tasks across 833 Android applications. contains 15,283 human demonstrations covering 14,548 unique tasks across 833 Android applications. This dataset is divided into a training set, an in-distribution (IDD) test set, a task-level OOD test set, a category-level OOD test set, and an app-level OOD test set. Each sample consists of a user goal, an observation, and a ground-truth action.
Metrics. For the online datasets AndroidWorld and AndroidLab, where the App agent is required to complete user-specified goals within an Android emulator, we evaluate performance using the task-level success rate (TSR), which is defined as the number of successfully completed tasks divided by the total number of tasks. A task is considered successful if the agent completes the goal within the dataset’s predefined maximum number of steps, and a failure otherwise. For the offline dataset AndroidControl, where the agent uses a user goal, an observation, and an action history as inputs to generate an action matching a predefined target, we use the step-level success rate (SSR), which is defined as the number of correctly predicted steps divided by the total number of steps. A step is successful if the generated action exactly matches the ground truth action; otherwise, it is a failure.
Implementation details. For the semantic evaluator, we use Gemini-Flash-2.0 gemini2024 as the LLM to predict the change that an action will bring to the UI state. The initial learning rate was set to 5e-5, with a linear decay schedule. Batch size was set to 1 and trained for 3 epochs. Mixed precision training (fp16) was enabled to reduce memory consumption and accelerate training. The training dataset includes 2K human-annotated step-wise samples from apps in the AndroidWorld emulator environment and 70K samples from the AndroidControl benchmark. The 2K human-generated samples are curated to ensure no overlap with the OOD test set in AndroidControl. We adopt a modified version of the prompt used in T3A, a leading prompt-driven App agent that operates entirely on textual inputs. While T3A uses two separate prompts, one for the action agent and another for the action summary agent that generates action summaries for use as history, our design uses a single prompt that instructs the agent to output both the action for app interaction and a concise textual description of the action, which is reused as the action history. In contrast to T3A, which requires the agent to provide a rationale for each action, our agent only generates the action and its description. This results in a fully text-based agent that is computationally efficient and does not rely on additional modules for generating summaries. We refer to the agent that uses our prompt as the ASL Agent.
Table 1: Comparison of different agents in terms of success rate in the AndroidWorld environment.
@@ 4.1 Experimental Setup
ASL-Agent Qwen2.5-7B-Instruct Text 5.6 0.0 0.0 2.5
Fine-Tuned ASL-Agent Qwen2.5-7B-Instruct-SFT Text 72.2 26.8 3.13 42.5
@@
4.2 Online Comparison
In this experiment, we evaluated model performance in real-time interactive environments using the AndroidWorld rawles2024androidworld and AndroidLab xu2024androidlab datasets. This setup is well-suited for testing our core motivation: that models should go beyond syntactic patterns and learn the semantics of actions—that is, their actual outcomes. In online testing, instead of evaluating each step in isolation, it assesses success based on whether the user goal is reached. Specifically, AndroidWorld provides an emulator that mimics real-world conditions by automatically executing actions and employs a hand-crafted rule-based system to verify task completion by checking system status. For AndroidLab, we use 30 provided tasks, excluding those involving question answering. We evaluate AndroidLab by executing actions within the AndroidWorld emulator and using an LLM-based evaluator to determine whether each task is completed successfully.
Table 1 presents the results on AndroidWorld, reporting success rates across tasks of varying difficulty—categorised as easy, medium, and hard—as well as the overall success rate. As shown in the table, our method outperforms the baseline by a significant margin of 3.8% in overall success rate. The most substantial improvements are observed in medium and hard tasks. Notably, our model maintains strong performance across all difficulty levels, especially where reasoning and semantic understanding of actions are more challenging. We further evaluate our method on the AndroidLab benchmark, with results presented in Table 2. The results are consistent to the results on AndroidWorld. Specifically, our model outperforms the pure syntax learning baseline Qwen2.5-7B-Instruct-SFT, and achieves the highest success rate. Together, these results highlight the generalisation capability of our model in real-world mobile UI environments and underscore the effectiveness of our ASL framework.
@@ 4.2 Online Comparison
Fine-Tuned ASL-Agent Qwen-2.5-7B-Instruct-SFT Text 33.3
@@
4.3 Offline Comparison
In this experiment, we evaluated model performance on the offline benchmark AndroidControl android_control . Specifically, we compared predicted actions against ground truth annotations. For action types, we assess whether the predicted action type string exactly matches the ground truth. For actions that involve coordinates, such as click and long press, we follow the evaluation protocol rawles2024androidinthewild to determine whether the ground truth coordinate falls within the bounding box of the predicted UI element.
AndroidControl provides four evaluation splits: IDD (In-Distribution Data), where all UI states and actions are sampled from the same distribution as the training data; Task-Unseen, where the task instructions are not observed during training; Cat-Unseen, where the UI categories are unseen; and App-Unseen, which involves entirely unseen applications. As shown in Table 3, we collect and evaluate both prompt-driven and models fine-tuned on AndroidControl.
Our model achieves the highest performance across all four splits. Under the IDD setting, it outperforms the strongest baseline with a significant margin. Moreover, it demonstrates strong generalisation to out-of-distribution scenarios. On the Task-Unseen split, our model surpasses the previous best. Similar improvements are observed on the Cat-Unseen and App-Unseen splits, highlighting the model’s robustness to new tasks, categories, and applications.
Table 3: Comparison across agents of step accuracy in the four splits of the AndroidControl test set.
@@ 4.3 Offline Comparison
Prompt-Driven SeeAct GPT-4o Image + Text 31.5 30.7 30.6 30.9
T3A GPT-4o Text 56.1 55.8 56.5 54.2
M3A GPT-4o Image + Text 60.8 59.3 60.8 60.4
ASL-Agent Qwen2.5-7B-Instruct Text 26.1 26.5 26.5 26.9
Fine-Tuned ASL-Agent Qwen2.5-7B-Instruct-SFT Text 67.4 61.2 59.7 59.7
@@ 4.4 Ablation Study
In this section, we study two key components of our ASL framework: the loss function and the reward model used during training. Ablation studies are conducted using the Qwen2.5-7B-Instruct model, trained on 3K samples from the AndroidControl training set and evaluated on 1K samples from the AndroidControl OOD test set. Results for the loss function ablation are presented in Table 5, where $\ell^{\text{SFT}}_{\theta}$ denotes the standard SFT loss used in the syntax learning paradigm, $\lambda{i},\ell^{\text{SFT}}_{\theta}$ represents the weighted SFT loss in our framework, and $\ell^{\text{ASL}}_{\theta}$ refers to the full loss function used in ASL. The results demonstrate that $\ell^{\text{ASL}}_{\theta}$ leads to the highest performance, highlighting its effectiveness in enabling the App agent to learn action semantics and improving robustness to OOD scenarios.
Table 5 reports the ablation results on the reward model. In this table, LLM denotes a baseline that directly uses Gemini-2.0-Flash to compute the semantics reward, while SEE refers to our proposed semantics estimator. The superior performance of SEE confirms its benefit in providing more informative and stable reward signals during training. These findings validate the effectiveness of both the loss function and the SEE reward model as core components of the ASL framework.

### v2

@@ 3.3 SEE: Semantic Estimator
where $o_{t}$ is the textual description of the UI state $s_{t}$, $a_{t}$ is the textual description of an action, and $\Delta\hat{o}_{t}$ is the description of the impact of the action $a_{t}$ on the UI state $s_{t}$. In practice, we implement this world model based on the proprietary LLM API by leveraging the LLM’s strong capabilities in language understanding and world knowledge to simulate how the environment would evolve in response to the action. The prompt of our world model can be found in the supplementary material.
Semantic calculator. With the state transitions generated by the world model, SEE employs a semantic calculator to generate the semantic reward for our ASL framework-generated transitions in two steps. Firstly, the semantic calculator uses BERT [7] to project the generated textual description for state transitions into a feature space. This process can be formulated as:
@@
This similarity score is used in our ASL framework as the semantic reward. A higher similarity score indicates a higher probability that a semantic equivalence exists between the action generated by the App agent and the ground truth action, providing a reward signal that guides the training of the agent towards semantically meaningful actions, even when the syntax may differ. This semantic estimator introduces minimal training-time overhead, as the world model is inference-only and is queried just once for each state–action pair.
Table 1: Comparison of different agents in terms of success rate in the AndroidWorld environment. T3A* refers to our prompt adapted from T3A, presented in the supplementary material.
@@
T3A* Qwen2.5-7B-Instruct Text 5.6 0.0 0.0 2.5
Fine-Tuned T3A* Qwen2.5-7B-Instruct-SFT Text 72.2 26.8 3.13 42.5
@@
3.4 Theoretical Analysis
In this section, we demonstrate that enabling the App agent to learn action semantics by minimising the semantic loss function defined in Eq. (4) leads to an agent that is more robust to the OOD problems compared to an agent that purely optimises the learning objective formulated in Eq. (1). Mathematically, we present Theorem 3.2 with the proof in the supplementary material.
Theorem 3.2.
4 Experiments
In this section, we first introduce our experimental setup. We then demonstrate the practical value of the proposed ASL framework on two online smartphone operation Apps benchmarks. To evaluate the robustness of the agent fine-tuned by our ASL framework, we also use an offline smartphone operation benchmark. Additionally, we interpret our semantic similarity score as a step-wise reward within an online RL fine-tuning pipeline, and further demonstrate its effectiveness. Finally, we conduct an ablation study to study the contribution of components within the ASL framework.
4.1 Experimental Setup
Datasets. We conducted our experiments on five benchmark datasets: AndroidLab [35], AndroidWorld [27], AndroidControl [16], Android-in-the-Wild (AitW) [28], and WebArena-Lite [43, 44]. These datasets encompass a diverse range of mobile UI settings and user interaction tasks, enabling a comprehensive evaluation of our model’s generalisation and understanding capabilities.
Metrics. For the online datasets such as AndroidWorld, performance is measured by the task-level success rate (TSR), i.e., the ratio of completed tasks to total tasks, where a task is successful if finished within the dataset’s step limit. For the offline dataset AndroidControl, we use the step-level success rate (SSR), i.e., the ratio of correctly predicted steps to total steps, where a step is successful if the generated action matches the ground truth
Implementation details. For the semantic evaluator, we use Gemini-Flash-2.0 [11] as the LLM to predict the change that an action will bring to the UI state. Batch size was set to 1 and trained for 3 epochs. The training dataset includes 2K human-annotated step-wise samples from apps in the AndroidWorld emulator environment and 70K samples from the AndroidControl benchmark. The 70K AndroidControl samples are collected to ensure no overlap with the test set in AndroidControl. Low-level instructions from AndroidControl are not used. T3A* refers to our prompt adapted from T3A, detailed in the supplementary material. Qwen-2.5-7B-Instruct-ASL refers to the model fine-tuned with our ASL loss, as defined in Eq. (5).
We report experiments on the AndroidWorld, AndroidLab, and AndroidControl datasets using ASL fine-tuned Qwen-2.5-7B. For the AitW tasks and WebArena-Lite benchmark, we follow Eq. (LABEL:eq:digirl_see_loss) and integrate our SEE module into their RL fine-tuning framework.
@@ 4.1 Experimental Setup
Fine-Tuned T3A* Qwen-2.5-7B-Instruct-SFT Text 33.3
@@
Table 3: Comparison across agents of step success rate in the four splits of AndroidControl.
@@
Prompt-Driven SeeAct Gemini-2.5-Flash Image + Text 36.7 35.9 33.7 34.4
GPT-4o 31.5 30.7 30.6 30.9
M3A Gemini-2.5-Flash 49.4 47.9 47.9 47.9
GPT-4o 60.8 59.3 60.8 60.4
T3A Gemini-2.5-Flash Text 49.1 45.4 44.4 45.0
GPT-4o 56.1 55.8 56.5 54.2
T3A* Gemini-2.5-Flash 46.8 45.3 44.7 43.5
Qwen2.5-7B-Instruct 26.1 27.2 26.7 26.3
Fine-Tuned T3A* Qwen2.5-7B-Instruct-SFT Text 67.4 61.2 59.7 59.7
@@
Table 4: Evaluation of our SEE module on the AitW General and Web Shopping tasks.
Task AppAgent CogAgent AutoUI Filtered BC Filtered BC-SEE DigiRL DigiRL-SEE
General Train 14.6% 25.0% 12.5% 53.5% 60.4% (+6.9%) 64.9% 70.8%(+5.9%)
Test 16.7% 25.0% 14.6% 62.5% 63.5% (+1.0%) 67.7% 71.9%(+4.2%)
Web Shopping Train 5.2% 31.3% 14.6% 53.6% 55.9%(+2.3%) 55.3% 57.3%(+2.0%)
Test 8.3% 38.5% 17.7% 54.2% 56.3% (+2.1%) 41.3% 55.2% (+13.9%)
4.2 Online Dataset Comparison
In this experiment, we evaluated model performance in real-time interactive environments using the AndroidWorld [27] and AndroidLab [35] datasets. This setup is well-suited for testing our core motivation: that models should go beyond syntactic patterns and learn the semantics of actions—that is, their actual outcomes. Specifically, AndroidWorld provides an emulator that mimics real-world conditions by automatically executing actions and employs a hand-crafted rule-based system to verify task completion by checking system status. For AndroidLab, we use 30 provided tasks, excluding those involving question answering. We evaluate AndroidLab by executing actions within the AndroidWorld emulator and using an LLM-based evaluator to determine whether each task is completed successfully. Table 1 presents the results on AndroidWorld, reporting success rates across tasks of varying difficulty—categorised as easy, medium, and hard—as well as the overall success rate. As shown in the table, our method outperforms the baseline by a significant margin of 3.8% in overall success rate. The most substantial improvements are observed in medium and hard tasks. We further evaluate our method on the AndroidLab benchmark, with superior results presented in Table 2. Together, these results highlight the generalisation capability of our model in real-world mobile UI environments and underscore the effectiveness of our ASL framework.
Table 5: Evaluation of our SEE module on the WebArena-Lite benchmark.
SFT Filtered BC Filtered BC-SEE DigiRL DigiRL-SEE
Task SSR 20.0% 23.0% 25.5% (+2.5%) 30.3% 32.1% (+1.8%)
4.3 Offline Dataset Comparison
In this experiment, we evaluated model performance on the offline benchmark AndroidControl [16]. Specifically, we compared predicted actions against ground truth annotations. For action types, we assess whether the predicted action type string exactly matches the ground truth. For actions that involve coordinates, such as click and long press, we follow the evaluation protocol [28] to determine whether the ground truth coordinate falls within the bounding box of the predicted UI element. As shown in Table 3, we collect and evaluate both prompt-driven and models fine-tuned on AndroidControl. Our model achieves the highest performance across all four splits, highlighting the model’s robustness to new tasks, categories, and applications.
4.4 RL Fine-tuning with SEE
To further validate the effectiveness, we evaluate our SEE module with different RL training pipelines on the AitW tasks [28] and WebArena-Lite benchmark [43, 44]. Following prior work [2, 34], we train on the full AitW training task split and evaluate on the first 96 tasks from both train and test splits of the General and Web Shopping categories, with success judged by Gemini [6]. For WebArena-Lite, all methods are trained and evaluated on all 165 tasks, with success assessed by a pretrained outcome reward model. Additional details are provided in the supplementary material. As shown in Table 4 and Table 5, both baselines consistently benefit from our SEE module. On AitW, Filtered BC-SEE outperforms Filtered BC by 6.9% on General train, while DigiRL-SEE achieves the best overall performance, surpassing DigiRL by 5.9% and 4.2% on General train and test. On WebArena-Lite, Filtered BC-SEE improves by 2.5%, and DigiRL-SEE further outperforms DigiRL by 1.8%. These results indicate that our SEE module provides denser training signals, alleviates the sparse reward issue, and enhances generalisation across mobile app and web control tasks.
@@ 4.5 Ablation Study
In this section, we study two key components of our ASL framework: the loss function and the reward model used during training. Ablation studies are conducted using the Qwen2.5-7B-Instruct model, trained on 3K samples from the AndroidControl training set and evaluated on 1K samples from the AndroidControl OOD test set. Results for the loss function ablation are presented in Table 6, where $\ell^{\text{SFT}}_{\theta}$ denotes the standard SFT loss used in the syntax learning paradigm, $\lambda_{i}\ell^{\text{SFT}}_{\theta}$ represents the weighted SFT loss in our framework, and $\ell^{\text{ASL}}_{\theta}$ refers to the full loss function used in ASL. The results demonstrate that $\ell^{\text{ASL}}_{\theta}$ leads to the highest performance, highlighting its effectiveness in enabling the App agent to learn action semantics and improving robustness to OOD scenarios.
Table 7 reports the ablation results on the reward model. In this table, ‘Gemini-2.0-Flash’ denotes a baseline that directly uses Gemini-2.0-Flash to compute the semantics reward, while SEE refers to our proposed semantics estimator. The superior performance of SEE confirms its benefit in providing more informative and stable reward signals during training. These findings validate the effectiveness of both the loss function and the SEE reward model as core components of the ASL framework.